# Bifurcation sequences in the secular 3D planetary 3-Body problem: a geometric approach

### v2

@@ 4.1 Integrable Hamiltonian $\mathcal{H}_{int}$
As a first example of application of the geometric method exposed to Section 3, we consider the integrable Hamiltonian $\mathcal{H}_{int}$ (Eq. (5)). Up to terms of second order in the variables $\sigma_{i}$ (i.e. of fourth order in the eccentricities), apart from constants, we find the approximate formula
$\mathcal{H}_{int}\approx A\sigma_{1}^{2}+C\sigma_{3}^{2}+B\sigma_{1}\sigma_{3}+D(\sigma_{0})\sigma_{1}+E(\sigma_{0})\sigma_{3}+F(\sigma_{0})$
where, for the parameters (masses, semi-major axes, and AMD value) as in the $\upsilon$-Andromedæ system (see [13]) the coefficients read:
$A\!=0.00212824,$ $C\!=0.00186469,$
$B\!=-0.00186482,$ $D(\sigma_{0})\!=0.000165361-0.0159745\,\sigma_{0},$
@@
The coefficients $A,B,C$ satisfy $B^{2}<4AC\,$, hence the quadratic form $A\sigma_{1}^{2}+C\sigma_{3}^{2}+B\sigma_{1}\sigma_{3}$ yields ellipses. Thus, for any permissible value $\mathcal{E}$, the curve $\mathcal{C}_{\sigma_{0},\,\mathcal{E}}^{(\sigma_{2}=0)}$ is ellipse-like.
@@
Figure 6: The first and third rows show the contour plots $\mathcal{P}_{\mathcal{H}_{int}}(\sigma_{0};\mathrm{AMD})$ (same as in Fig. 2), where the position of the fixed points (indicated by colored points) has been computed by the tangency method (see text). The second and fourth rows show the intersections of the spheres $\mathcal{S}_{\sigma_{0}}$ (blue curve $\mathcal{S}_{\sigma_{0}}^{(\sigma_{2}=0)}$) with the energy surfaces $\mathcal{C}_{\sigma_{0},\mathcal{E}}$(red dashed curves $\mathcal{C}_{\sigma_{0},\mathcal{E}}^{(\sigma_{2}=0)}$) with the plane $(\sigma_{1},\sigma_{3})$ for $\sigma_{2}=0$, for different values of the energy $\mathcal{E}$ and values of $\sigma_{0}$ as indicated in each panel.
@@
Figure 6 shows the phase portraits $\mathcal{P}_{\mathcal{H}_{int}}(\sigma_{0})$ in the above model. Altering the value of $\sigma_{0}$ as indicated in each panel, we obtain fixed points of both classes CP1 and CP2 in the phase portraits (first and third row). These are compared with the fixed points found through the curves $\mathcal{S}_{\sigma_{0}}^{(\sigma_{2}=0)}$ and $\mathcal{C}_{\sigma_{0},\mathcal{E}}^{(\sigma_{2}=0)}$. As $\sigma_{0}$ decreases, implying a smaller value of the total angular momentum, the general level of mutual inclination of the orbits increases (from top left to bottom right in Fig. 6). The top left panel shows the starting point of the observed bifurcation sequence, where we have the two fixed points A and B corresponding to the basic ACR states. These are tangency points of $\mathcal{S}_{\sigma_{0}}$ with $\mathcal{C}_{\sigma_{0},\,\mathcal{E}}$ at the values $\sigma_{0}=\sigma_{0_{max}}(=1.62\cdot 10^{-2}$) and the energies $\mathcal{E}=\mathcal{E}_{min}$ for the mode A, and $\mathcal{E}=\mathcal{E}_{2,3}$ for the mode B (cf. with caption in Fig. 3). Both A and B are points of outer tangency of the curves $\mathcal{S}_{\sigma_{0}}^{(\sigma_{2}=0)}$ and $\mathcal{C}_{\sigma_{0},\mathcal{E}}^{(\sigma_{2}=0)}$, hence linearly stable. It is easy to verify that $\psi^{(A)}=0$ and $\psi^{(B)}=\pi$, leading to $\varpi_{3}=\varpi_{2}+\pi$ (perihelia anti-aligned) and $\varpi_{3}=\varpi_{2}$ (perihelia aligned). Hence, the associated orbits yield two coplanar ellipses with, respectively, anti-aligned and aligned pericenters precessing with the same frequency.
@@
Decreasing the value of $\sigma_{0}$, at the second frame on the top of Fig. 6, a saddle-node bifurcation takes place, giving rise to two new fixed points, i.e., the periodic orbits $P_{1}$ and $P_{2}$. $P_{1}$ corresponds to an inner tangency of the curves $\mathcal{S}_{\sigma_{0}}^{(\sigma_{2}=0)}$ with $\mathcal{C}_{\sigma_{0},\mathcal{E}}^{(\sigma_{2}=0)}$ (unstable), while $P_{2}$ corresponds to an outer tangency (stable) (see Fig. 4). Besides these orbits, we still have present in the phase portraits the stable fixed points A and B (outer tangencies) of the ACR states. However, the ellipse yielding the outer tangency of the fixed point A gradually shrinks in size as $\sigma_{0}$ decreases. Then, at a critical value of $\sigma_{0}=\sigma_{0}^{(A)}$ the ellipse shrinks to a point, while, for still lower values, as in the third top panel of Fig. 6, the tangent ellipse re-emerges, being now contained in the disc limited by $\mathcal{S}_{\sigma_{0}}^{(\sigma_{2}=0)}$, yielding now a fixed point A of inner tangency, hence unstable. At the critical value $\sigma_{0}^{(A)}$ we then have a pitchfork bifurcation, accompanied with the birth of two new stable fixed points ($F_{1}$ and $F_{2}$), not shown in the second and fourth rows of Fig. 6 since they are of the type CP2, i.e., out of the plane $\sigma_{2}=0$. For values of $\sigma_{0}$ smaller than $\sigma_{0}^{(A)}$, there are surfaces $\mathcal{C}_{\sigma_{0},\,\mathcal{E}}$ of the form of the elliptic cylinders, as in the left panel of Fig. 5, which intersect transversely with the sphere $\mathcal{S}_{\sigma_{0}}$. The intersection yield invariant curves forming the islands of stability around the fixed points $F_{1}$ and $F_{2}$. The latter correspond to the value of the energy $\mathcal{E}^{(F)}$ at which the elliptic cylinders collide to straight lines (middle panel in Fig. 5). In the third, fourth and fifth panel we find elliptic cylinders for energies $\mathcal{E}>\mathcal{E}^{(F)}$. However, decreasing further $\sigma_{0}$, the fixed points $F_{1}$, $F_{2}$ eventually collide with $P_{1}$ (sixth panel in Fig. 6), and this collision terminates the $F$-family of periodic orbits by an inverse pitchfork bifurcation which renders the point $P_{1}$ stable. Note that this process is equivalent to the one of the birth of the F-family, since, at the point of the collision the tangency corresponding to the point $P_{1}$ turns from inner to outer, i.e., the point $P_{1}$ turns from unstable to stable. Finally, decreasing $\sigma_{0}$ still further, the points $P_{2}$ and A similarly disappear by an inverse saddle-node bifurcation. Hence, the only surviving periodic orbits for small $\sigma_{0}$ are the stable orbits $P_{1}$ and B.
@@ 4.2 Secular normal form $\mathcal{H}_{int}^{(1)}$
To compute the normal form, we make use of the ‘book-keeping method’ discussed in [4]. Starting from the Hamiltonian of Eq. (13) written in the canonical variables of Eq. (3.1)
$\mathcal{H}_{sec}(\psi,\varphi,\Gamma,J)=\mathcal{H}_{int}(\psi,\Gamma;J)+\mathcal{H}_{1,space}(\psi,\varphi,\Gamma,J)\,$
$\mathcal{H}_{int}$ is decomposed as
$\mathcal{H}_{int}(\psi,\Gamma;J)=Z_{0}(J)+\lambda Z_{0,1}(\psi,\Gamma;J)\,.$
@@
$\mathcal{H}^{(0)}=\mathcal{H}_{sec}(\psi,\varphi,\Gamma,J)=Z_{0}(J)+\lambda Z_{0,1}(\psi,\Gamma;J)+\lambda\mathcal{H}_{1,space}(\psi,\varphi,\Gamma,J)\,.$ (43)
All the terms contained in $\mathcal{H}_{1,space}$ depend on the angle $\varphi$. We then define the Lie generating function $\chi$ as the solution of the following homological equation:
@@
To solve the homological equation, we first re-express the Hamiltonian in the variables $(w_{2},w_{3},W_{2},W_{3})$, introduced in Eq. (11), rather than $(\psi,\varphi,\Gamma,J)$. Then
$\mathcal{H}_{sec}(w_{2},w_{3},W_{2},W_{3})=\mathcal{H}_{int}(w_{2}-w_{3},W_{2},W_{3})+\mathcal{H}_{1,space}(w_{2},w_{3},W_{2},W_{3})\,$
where
$\mathcal{H}_{int}=Z_{0}(W_{2},W_{3})+\lambda Z_{0,1}(w_{2}-w_{3},W_{2},W_{3}),$
with
$Z_{0}(W_{2},W_{3})=c+aW_{2}+bW_{3}\leavevmode\nobreak\ \leavevmode\nobreak\ .$ (45)
We then obtain
where $\mathcal{H}_{1,space}(w_{2},w_{3},W_{2},W_{3})$ contains trigonometric terms depending on linear combinations of the angles $k_{1}w_{2}+k_{2}w_{3}$ with $k_{1}+k_{2}\neq 0$. Thus, writing the Taylor-Fourier expansion of $\mathcal{H}_{1,space}$ as
@@
[…]
@@
$\mathcal{H}^{(1)}_{int}=\mathcal{H}_{int}+\Bigg{[}\left\langle\frac{\lambda}{2}\{\mathcal{H}_{1,space},\chi\}\right\rangle_{\varphi}\Bigg{]}_{\leq N_{bk}}\leavevmode\nobreak\ \leavevmode\nobreak\ ,$ (50)
where the operation $[\cdot]_{N_{bk}}$ means to truncate the enclosed expression at the maximum adopted order of expansion in the eccentricities and inclinations ($N_{bk}=12$ in our examples below). The integrable Hamiltonian model $\mathcal{H}^{(1)}_{int}$ is hereafter referred to as the first order secular normal form (or simply the ‘secular normal form’).
Phase portraits under the secular normal form $\mathcal{H}_{int}^{(1)}$ can be obtained using the change of variables (3.1) and (16) and computing Poincaré surfaces of section $\mathcal{PS}_{\mathcal{H}_{int}^{(1)}}(\mathcal{E};\mathrm{AMD})$ at a fixed level of energy $\mathcal{E}$ through
[…]
@@
Figure 7: Poincaré surfaces of section $\mathcal{PS}_{\mathcal{H}_{int}^{(1)}}(\mathcal{E};\mathrm{AMD})$ in the plane $(X_{2},Y_{2})$ with $\mathrm{AMD}$ fixed. The surfaces of section are computed by a numerical integration of trajectories of the normalized integrable Hamiltonian ${\mathcal{H}}^{(1)}_{int}$ (Eq. (50)) truncated at multipolar degree $N_{P}=5$, order $N_{bk}=12$ in the eccentricities, and energies (from top to bottom, from left to right) $\mathcal{E}=-5.82\cdot 10^{-5},-4.3\cdot 10^{-5},-3.59\cdot 10^{-5},-3.34\cdot 10^{-5},-3.03\cdot 10^{-5},-2.68\cdot 10^{-5}$,$-2.51\cdot 10^{-5},-2.42\cdot 10^{-5},-2.12\cdot 10^{-5},-1.99\cdot 10^{-5},-1.44\cdot 10^{-5},-3.3\cdot 10^{-6}$. In this case the energies for which we can have co-planar orbits range between $\mathcal{E}_{min}=-1.71\cdot 10^{-4}$ and $\mathcal{E}_{2,3}=-5.82\cdot 10^{-5}\,$. In the plot we consider the phase-portraits for energies $\mathcal{E}\geq\mathcal{E}_{2,3}$, for which no-coplanar orbits exist.
@@
Figure 7 shows the surfaces of section for different (increasing, from left to right, from top to bottom) levels of energy. By visual inspection, the sequence of bifurcations produced by the secular normal form $\mathcal{H}_{int}^{(1)}$ differ considerably from those of the integrable hamiltonian $\mathcal{H}_{int}$ before the normalization (top row of Fig. 2). The origin of the difference becomes clear by the analysis through Hopf variables as discussed below. On the other hand, comparison with the phase portraits of the full secular Hamiltonian $\mathcal{H}_{sec}$ (Fig. 1) shows that the secular normal form reproduces qualitatively the same sequence of bifurcations as in the complete model, except for the last bifurcation (transition from the fourth to the fifth panel in Fig. 1), connected to the birth of the Lidov-Kozai periodic orbits. Such orbits are associated with the $e_{2}^{2}\cos(2w_{2})$ or $e_{3}^{2}\cos(2w_{3})$ terms in the secular Hamiltonian, which are not present in the secular normal form.
@@
We now apply the geometric method of section 3 to the secular normal form $\mathcal{H}_{int}^{(1)}$ to recover the sequence of bifurcations observed in the regime of energies $\mathcal{E}_{C}\leq\mathcal{E}\leq\mathcal{E}_{C,2}$, where $\mathcal{E}_{C}$ is the energy of the saddle-node bifurcation of the orbits $P_{1}$ and $P_{2}$, and $\mathcal{E}_{C,2}$ is the energy of bifurcation of the Lidov-Kozai orbits. Similarly to Fig. 6, in the case of the secular normal form Fig. 8 yields the phase portraits of the secular normal form expressed in Hopf variables $\mathcal{H}_{int}^{(1)}(\sigma_{1},\sigma_{3};\sigma_{0})$. Each upper panel in Fig. 8 corresponds to a fixed value of $\sigma_{0}$ decreasing from left to right and from top to bottom, and the invariant curves in the same portrait correspond to different values of the energy $\mathcal{E}=\mathcal{H}_{int}^{(1)}$ within the allowed limits computed in the same way as before for the Hamiltonian $\mathcal{H}_{int}$. The lower panel below each phase portrait shows the intersections or tangencies of various curves $\mathcal{C}_{\sigma_{0},\,\mathcal{E}}^{(\sigma_{2}=0)}$ with the circle $\mathcal{S}_{\sigma_{0}}^{(\sigma_{2}=0)}$ for the corresponding value of $\sigma_{0}$.
@@
Comparing Figs. 6 and 8, we immediately notice that, for the same model parameters, the curves $\mathcal{S}_{\sigma_{0}}^{(\sigma_{2}=0)}$ are hyperbola-like in the case of the secular normal form, instead of ellipse-like, as in the Hamiltonian $\mathcal{H}_{int}$. Up to terms of second order in the variables $\sigma_{i}$, the secular normal form reads (apart from constants)
where, in our numerical example
$A^{(1)}\!=0.00912208,$ $C^{(1)}\!=-0.0518887,$
@@
Hence, $(B^{(1)})^{2}>4A^{(1)}C^{(1)}\,$, implying that the quadratic form $A^{(1)}\sigma_{1}^{2}+C^{(1)}\sigma_{3}^{2}+B^{(1)}\sigma_{1}\sigma_{3}$ yields now a family of hyperbolas instead of ellipses. This difference brings essential differences in the structure of the phase portraits of the secular normal form $\mathcal{H}_{int}^{(1)}$ compared to the integrable model $\mathcal{H}_{int}$.

### v3

@@ 4.1 Integrable Hamiltonian $\mathcal{H}_{int}$
As a first example of application of the geometric method exposed to Section 3, we consider the integrable Hamiltonian $\mathcal{H}_{int}$ (Eq. (5)). Apart from constants, the integrable Hamiltonian can be written as
$\mathcal{H}_{int}=A\sigma_{1}^{2}+C\sigma_{3}^{2}+B\sigma_{1}\sigma_{3}+D(\sigma_{0})\sigma_{1}+E(\sigma_{0})\sigma_{3}+F(\sigma_{0})+\mathcal{O}\left((\sigma_{i})^{2}\right)\,$, (52)
@@
with $i=0,\,\dots,3$, where the symbol $\mathcal{O}((\sigma_{i})^{2})$ correspond to fourth order in the eccentricities. The numerical value of the coefficients are given in Appendix A. For those values of the coefficients $A,B,C$, the condition $B^{2}<4AC\,$ is satisfied, hence the quadratic form $A\sigma_{1}^{2}+C\sigma_{3}^{2}+B\sigma_{1}\sigma_{3}$ yields ellipses. Thus, for any permissible value $\mathcal{E}$, the curve $\mathcal{C}_{\sigma_{0},\,\mathcal{E}}^{(\sigma_{2}=0)}$ is ellipse-like.
@@
Figure 7: The first and third rows show the contour plots $\mathcal{P}_{\mathcal{H}_{int}}(\sigma_{0};\mathrm{AMD})$ (same as in Fig. 3), where the position of the fixed points (indicated by colored points) has been computed by the tangency method explained in Section 3.3, i.e. computing, for each panel, the tangencies between the sphere $\mathcal{S}_{\sigma_{0}}$ of fixed radius $\sigma_{0}$ and $\mathcal{C}_{\sigma_{0},\mathcal{E}}.$ Computing these tangencies, we obtain critical points of the first kind (i.e. A, B, $P_{1}$, $P_{2}$) satisfying the condition given by Eq. (36), as well as critical points of the second kind (the points $F_{1}$ and $F_{2}$) satisfying (37). The second and fourth rows show the intersections of the spheres $\mathcal{S}_{\sigma_{0}}$ (blue curve $\mathcal{S}_{\sigma_{0}}^{(\sigma_{2}=0)}$) with the energy surfaces $\mathcal{C}_{\sigma_{0},\mathcal{E}}$(red dashed curves $\mathcal{C}_{\sigma_{0},\mathcal{E}}^{(\sigma_{2}=0)}$) in the plane $(\sigma_{1},\sigma_{3})$ for $\sigma_{2}=0$, for different values of the energy $\mathcal{E}$ and values of $\sigma_{0}$ as indicated in each panel.
@@
Figure 7 shows the phase portraits $\mathcal{P}_{\mathcal{H}_{int}}(\sigma_{0})$ in the above model. Altering the value of $\sigma_{0}$ as indicated in each panel, we obtain fixed points of both classes CPI and CPII in the phase portraits (first and third row). These are compared with the fixed points found through the curves $\mathcal{S}_{\sigma_{0}}^{(\sigma_{2}=0)}$ and $\mathcal{C}_{\sigma_{0},\mathcal{E}}^{(\sigma_{2}=0)}$. As $\sigma_{0}$ decreases, implying a smaller value of the total angular momentum, the general level of mutual inclination of the orbits increases (from top left to bottom right in Fig. 7). The top left panel shows the starting point of the observed bifurcation sequence, where we have the two fixed points A and B corresponding to the basic ACR states. These are tangency points of $\mathcal{S}_{\sigma_{0}}$ with $\mathcal{C}_{\sigma_{0},\,\mathcal{E}}$ at the values $\sigma_{0}=\sigma_{0_{max}}(=1.62\cdot 10^{-2}$) and the energies $\mathcal{E}=\mathcal{E}_{min}$ for the mode A, and $\mathcal{E}=\mathcal{E}_{2,3}$ for the mode B (cf. with caption in Fig. 2). Both A and B are points of outer tangency of the curves $\mathcal{S}_{\sigma_{0}}^{(\sigma_{2}=0)}$ and $\mathcal{C}_{\sigma_{0},\mathcal{E}}^{(\sigma_{2}=0)}$, hence linearly stable. It is easy to verify that $\psi^{(A)}=0$ and $\psi^{(B)}=\pi$, leading to $\varpi_{3}=\varpi_{2}+\pi$ (perihelia anti-aligned) and $\varpi_{3}=\varpi_{2}$ (perihelia aligned). Hence, the associated orbits yield two coplanar ellipses with, respectively, anti-aligned and aligned pericenters precessing with the same frequency.
@@
Decreasing the value of $\sigma_{0}$, at the second frame on the top of Fig. 7, a saddle-node bifurcation takes place, giving rise to two new fixed points, i.e., the periodic orbits $P_{1}$ and $P_{2}$. $P_{1}$ corresponds to an inner tangency of the curves $\mathcal{S}_{\sigma_{0}}^{(\sigma_{2}=0)}$ with $\mathcal{C}_{\sigma_{0},\mathcal{E}}^{(\sigma_{2}=0)}$ (unstable), while $P_{2}$ corresponds to an outer tangency (stable) (see Fig. 4). Besides these orbits, we still have present in the phase portraits the stable fixed points A and B (outer tangencies) of the ACR states. However, the ellipse yielding the outer tangency of the fixed point A gradually shrinks in size as $\sigma_{0}$ decreases. Then, at a critical value of $\sigma_{0}=\sigma_{0}^{(A)}$ the ellipse shrinks to a point, while, for still lower values, as in the third top panel of Fig. 7, the tangent ellipse re-emerges, being now contained in the disc limited by $\mathcal{S}_{\sigma_{0}}^{(\sigma_{2}=0)}$, yielding now a fixed point A of inner tangency, hence unstable. At the critical value $\sigma_{0}^{(A)}$ we then have a pitchfork bifurcation, accompanied with the birth of two new stable fixed points ($F_{1}$ and $F_{2}$), not shown in the second and fourth rows of Fig. 7 since they are of the type CPII, i.e., out of the plane $\sigma_{2}=0$. For values of $\sigma_{0}$ smaller than $\sigma_{0}^{(A)}$, there are surfaces $\mathcal{C}_{\sigma_{0},\,\mathcal{E}}$ of the form of the elliptic cylinders, as in the top left panel of Fig. 5, which intersect transversely with the sphere $\mathcal{S}_{\sigma_{0}}$. The intersection yield invariant curves forming the islands of stability around the fixed points $F_{1}$ and $F_{2}$. The latter correspond to the value of the energy $\mathcal{E}^{(F)}$ at which the elliptic cylinders collide to straight lines ( top right panel in Fig. 5). In the third, fourth and fifth panel we find elliptic cylinders for energies $\mathcal{E}>\mathcal{E}^{(F)}$. However, decreasing further $\sigma_{0}$, the fixed points $F_{1}$, $F_{2}$ eventually collide with $P_{1}$ (sixth panel in Fig. 7), and this collision terminates the $F$-family of periodic orbits by an inverse pitchfork bifurcation which renders the point $P_{1}$ stable. Note that this process is equivalent to the one of the birth of the F-family, since, at the point of the collision the tangency corresponding to the point $P_{1}$ turns from inner to outer, i.e., the point $P_{1}$ turns from unstable to stable. Finally, decreasing $\sigma_{0}$ still further, the points $P_{2}$ and A similarly disappear by an inverse saddle-node bifurcation. Hence, the only surviving periodic orbits for small $\sigma_{0}$ are the stable orbits $P_{1}$ and B.
@@ 4.2 Secular normal form $\mathcal{H}_{int}^{(1)}$
To compute the normal form, we make use of the ‘book-keeping method’ discussed in [4]. Starting from the Hamiltonian of Eq. (17)
$\mathcal{H}_{sec}(w_{2},w_{3},W_{2},W_{3})=\mathcal{H}_{int}(w_{2}-w_{3},W_{2},W_{3})+\mathcal{H}_{1,space}(w_{2},w_{3},W_{2},W_{3})\,$,
the integrable part of the Hamiltonian $\mathcal{H}_{int}$ is decomposed as
$\mathcal{H}_{int}=Z_{0}(W_{2},W_{3})+\lambda Z_{0,1}(w_{2}-w_{3},W_{2},W_{3}),$
@@
where $\mathcal{H}_{1,space}(w_{2},w_{3},W_{2},W_{3})$ contains trigonometric terms depending on linear combinations of the angles $k_{1}w_{2}+k_{2}w_{3}$ with $k_{1}+k_{2}\neq 0$. Using angles $(\psi,\varphi)$ (see Eq. (26)) this would have implied that all the terms contained in $\mathcal{H}_{1,space}$ depend on the angle $\varphi$. We then define the Lie generating function $\chi$ as the solution of the following homological equation:
@@
Thus, writing the Taylor-Fourier expansion of $\mathcal{H}_{1,space}$ as
@@
Note that the generating function $\chi$ can be defined for all systems with masses $m_{j}$ and semi-major axes $a_{j}$, $j=2,3$ such that we are not close to a secular resonance, i.e., a commensurability of the form a(n+k)=b(nk)a(n+k)=b(n-k). For any finite truncation of the Hamiltonian, $\mathcal{H}_{1,space}$ contains harmonics of the form $\cos((n+k)w_{2}-(n-k)w_{3})$ with a rather low order |nk|+|n+k||n-k|+|n+k|, which is anyway limited from above, hence the measure of the parameters $(a_{j},m_{j})$ leading to secular (near-)resonance is small.
@@
where $\exp L_{\chi}\cdot=\sum_{j\geq 0}\left(L_{\chi}^{j}\cdot\right)/j!$ is the Lie series operator, $L_{\chi}\cdot=\{\cdot,\chi\}$ is the Lie derivative with respect to the dynamical function $\chi$ and $\{\cdot,\cdot\}$ denotes the Poisson bracket. Expanding Eq. (56) in powers of the book-keeping symbol $\lambda$, we find
$\mathcal{H}^{(1)}$ $=\underbrace{Z_{0}+\overbracket{\lambda Z_{0,1}}^{\mathcal{O}(\lambda)}}_{\mathcal{H}_{int}}+\underbrace{\overbracket{\lambda\mathcal{H}_{1,space}}^{\mathcal{O}(\lambda)}+\overbracket{\{Z_{0},\chi\}}^{\mathcal{O}(\lambda)}}_{=\,0}+\overbracket{\lambda\{Z_{0,1},\chi\}}^{\mathcal{O}(\lambda^{2})}+\underbrace{\overbracket{\lambda\{\mathcal{H}_{1,space},\chi\}}^{\mathcal{O}(\lambda^{2})}+\frac{1}{2}\overbracket{\{\{Z_{0},\chi\},\chi\}}^{\mathcal{O}(\lambda^{2})}}_{=\,\frac{\lambda}{2}\{\mathcal{H}_{1,space},\chi\}}+\mathcal{O}(\lambda^{3})$ (57)
@@
$\mathcal{H}^{(1)}_{int}=\mathcal{H}_{int}+\Bigg[\left\langle\frac{\lambda}{2}\{\mathcal{H}_{1,space},\chi\}\right\rangle_{\varphi}\Bigg]_{\leq N_{bk}}\penalty 10000\ \penalty 10000\ ,$ (59)
where the operation $\cdot]_{N_{bk}}$ means to truncate the enclosed expression at the maximum adopted order of expansion in the eccentricities and inclinations ($N_{bk}=12$ in our examples below). The integrable Hamiltonian model $\mathcal{H}^{(1)}_{int}$ is hereafter referred to as the first order secular normal form (or simply the ‘secular normal form’).
Phase portraits under the secular normal form $\mathcal{H}_{int}^{(1)}$ can be obtained using the change of variables (26) and (20) and computing Poincaré surfaces of section $\mathcal{PS}_{\mathcal{H}_{int}^{(1)}}(\mathcal{E};\mathrm{AMD})$ at a fixed level of energy $\mathcal{E}$ through
$\mathcal{PS}_{\mathcal{H}_{int}^{(1)}}(\mathcal{E};\mathrm{AMD})=\bigg\{$ $(X_{2},Y_{2},X_{3},Y_{3})\in\mathbb{R}^{4}:\penalty 10000\ \mathcal{H}^{(1)}_{int}(X_{2},Y_{2},X_{3},Y_{3}=0;\mathrm{AMD})=\mathcal{E}\,,\,Y_{3}=0\,$, (60)
@@
Figure 8: Poincaré surfaces of section $\mathcal{PS}_{\mathcal{H}_{int}^{(1)}}(\mathcal{E};\mathrm{AMD})$ in the plane $(X_{2},Y_{2})$ with $\mathrm{AMD}$ fixed. The surfaces of section are computed by a numerical integration of trajectories of the normalized integrable Hamiltonian ${\mathcal{H}}^{(1)}_{int}$ (Eq. (59)) truncated at multipolar degree $N_{P}=5$, order $N_{bk}=12$ in the eccentricities, and energies (from top to bottom, from left to right) $\mathcal{E}=-5.82\cdot 10^{-5},-4.3\cdot 10^{-5},-3.59\cdot 10^{-5},-3.34\cdot 10^{-5},-3.03\cdot 10^{-5},-2.68\cdot 10^{-5}$$2.51\cdot 10^{-5},-2.42\cdot 10^{-5},-2.12\cdot 10^{-5},-1.99\cdot 10^{-5},-1.44\cdot 10^{-5},-3.3\cdot 10^{-6}$. In this case the energies for which we can have co-planar orbits range between $\mathcal{E}_{min}=-1.71\cdot 10^{-4}$ and $\mathcal{E}_{2,3}=-5.82\cdot 10^{-5}\,$. In the plot we consider the phase-portraits for energies $\mathcal{E}\geq\mathcal{E}_{2,3}$, for which no-coplanar orbits exist.
@@
Figure 8 shows the surfaces of section for different (increasing, from left to right, from top to bottom) levels of energy. By visual inspection, the sequence of bifurcations produced by the secular normal form $\mathcal{H}_{int}^{(1)}$ differ considerably from those of the integrable Hamiltonian $\mathcal{H}_{int}$ before the normalization (top row of Fig. 3). The origin of the difference becomes clear by the analysis through Hopf variables as discussed below. On the other hand, comparison with the phase portraits of the full secular Hamiltonian $\mathcal{H}_{sec}$ (Fig. 1) shows that the secular normal form reproduces qualitatively the same sequence of bifurcations as in the complete model, except for the last bifurcation (transition from the fourth to the fifth panel in Fig. 1), connected to the birth of the Lidov-Kozai periodic orbits. Such orbits are associated with the $\mathrm{e}_{2}^{2}\cos(2w_{2})$ or $\mathrm{e}_{3}^{2}\cos(2w_{3})$ terms in the secular Hamiltonian, which are not present in the secular normal form.
@@
We now apply the geometric method of section 3 to the secular normal form $\mathcal{H}_{int}^{(1)}$ , given by Eq. (59), to recover the sequence of bifurcations observed in the regime of energies $\mathcal{E}_{C}\leq\mathcal{E}\leq\mathcal{E}_{C,2}$, where $\mathcal{E}_{C}$ is the energy of the saddle-node bifurcation of the orbits $P_{1}$ and $P_{2}$, and $\mathcal{E}_{C,2}$ is the energy of bifurcation of the Lidov-Kozai orbits (see Section 2.2). Similarly to Fig. 7, in the case of the secular normal form Fig. 9 yields the phase portraits of the secular normal form expressed in Hopf variables $\mathcal{H}_{int}^{(1)}(\sigma_{1},\sigma_{3};\sigma_{0})$. Each upper panel in Fig. 9 corresponds to a fixed value of $\sigma_{0}$ decreasing from left to right and from top to bottom, and the invariant curves in the same portrait correspond to different values of the energy $\mathcal{E}=\mathcal{H}_{int}^{(1)}$ within the allowed limits computed in the same way as before for the Hamiltonian $\mathcal{H}_{int}$ (i.e., as already done in Fig. 2 for $\mathcal{H}_{int}$, it is sufficient to fix a value of $\sigma_{0}$ and find the limits on the energies trough the computations of the tangencies between Eq. (33) and Eq. (34), with $\mathcal{Z}=\mathcal{H}_{int}^{(1)}$). The lower panel below each phase portrait shows the intersections or tangencies of various curves $\mathcal{C}_{\sigma_{0},\,\mathcal{E}}^{(\sigma_{2}=0)}$ with the circle $\mathcal{S}_{\sigma_{0}}^{(\sigma_{2}=0)}$ for the corresponding value of $\sigma_{0}$.
@@
Comparing Figs. 7 and 9, we immediately notice that, for the same model parameters, the curves $\mathcal{S}_{\sigma_{0}}^{(\sigma_{2}=0)}$ are hyperbola-like in the case of the secular normal form, instead of ellipse-like, as in the Hamiltonian $\mathcal{H}_{int}$.
Apart from constants, the secular normal form reads
$\mathcal{H}_{int}^{(1)}=A^{(1)}\sigma_{1}^{2}+C^{(1)}\sigma_{3}^{2}+B^{(1)}\sigma_{1}\sigma_{3}+D^{(1)}(\sigma_{0})\sigma_{1}+E^{(1)}(\sigma_{0})\sigma_{3}+F^{(1)}(\sigma_{0})+\mathcal{O}\left((\sigma_{i})^{2}\right)\,$, (61)
@@
with $i=0,\,\dots,3$. The numerical value of the coefficients are given in Appendix B. For those values of the coefficients $A^{(1)},B^{(1)},C^{(1)}$, the condition $\left(B^{(1)}\right)^{2}>4A^{(1)}C^{(1)}\,$ is satisfied, implying that the quadratic form $A^{(1)}\sigma_{1}^{2}+C^{(1)}\sigma_{3}^{2}+B^{(1)}\sigma_{1}\sigma_{3}$ yields now a family of hyperbolas instead of ellipses. This difference brings essential differences in the structure of the phase portraits of the secular normal form $\mathcal{H}_{int}^{(1)}$ compared to the integrable model $\mathcal{H}_{int}$.